In Hydrogen atom electron transition takes place between L shell and K shell then wavelength of light emitted is

Grade:11

3 Answers

Godfrey Classic Prince
12 years ago

Dear moses christoper,

If an electronic transition takes place from the L Shell to the K Shell i.e., n=2(L Shell) to n=1(K Shell) then the wavelength of the light emitted is given by the Rydberg's formula

\frac{1}{\lambda_{vac}} = R\left(\frac{1}{n_1^2}-\frac{1}{n_2^2}\right)

Where

\lambda_{vac} \! is the wavelength of electromagnetic radiation emitted in vacuum,
R\! is the Rydberg constant, approximately 1.097 * 107 m-1,
n_1\! and n_2\! are integers such that n_1 < n_2\!.

Therefore the wavelength according to your problem will be 91.13 nm

Hope this helped you immensely...

Regards,

AskIITians Expert

Godfrey Classic Prince

IIT-Madras

Please approve my answer if you liked it by clicking on "Yes" given below...!!Smile

Askiitians_Expert Yagyadutt
12 years ago

Dear Christopher !

 

for k n=1 and for L n=2

 

Energy at every level is equal to  -13.6/n^2

 

So at n=1 it is -13.6 eV

 

and at n=2 it is -3.4 ev

 

differnece in the energy at two levels is  = 10.2 ev which is definately a negative...means if an electron jumps from L to K it will release energy ..

 

So wavelength corresponding to 10.2 ev will be emitted ..

 

hc/lamda = 10.2 ev

12400/lamda = 10.2

 

or lamda = 12400/10.2   A   (calculate for the answer)

 

regards

 

Yagya

Mukkamala Viswateja
12 years ago

i think the wavelength is present in between the uv region.